OpenStudy (kingdex04):

The terminal side of angle THETA lies along the line L given by Y = Sqrt(3)x+1 = 1,X less than or equal to 0. how can i find the exct 6 trig functions from this, or atleast the x and y values

OpenStudy (loser66):

The question is unclear to me. What is Y =Sqrt(3)x+1=1 ???

OpenStudy (loser66):

y/x = tan theta =-sqrt 3 = tan (2pi/3) Hence theta = 2pi/3, now you can find cos (2pi/3) sin(2pi/3) cot (2pi/3) and 2 more.